# Phase driven exact and unconventional superradiance phase transition in non-Hermitian cascaded Rabi cavities

### v1

@@ Abstract
This work reports the phase driven symmetry breaking and exact and unconventional superradiance phase transition in the non-Hermitian cascaded Rabi cavities. The non-Hermiticity is introduced in the coupling phase (denoted by $\varphi$) between the atom and the optical field. The exactness refers to the fact that the superradiance phase boundary is obtained analytically and verified by the observables. The unconventionality is reflected in that when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, the phase boundary is uniquely determined by $\mathcal{J}=\frac{1}{2}$ (where $\mathcal{J}$ is the dimensionless cavity coupling strength) and is independent of the atom-optical field coupling strength $g$. For other $\varphi$, the phase boundary is determined by $\mathcal{J}$ and the dimensionless atom-optical field coupling strength $g$ together. Besides, we find that there are phase driven first-order and second-order superradiance phase transitions, and the quantum criticality for the second-order superradiance phase transition is studied. In addition, the experimental feasibility is discussed. This work will stimulate the studies of non-Hermitian superradiance quantum phase transitions and their experimental realizations, as well as the underlying universality class of phase transitions.
@@ I Introduction
The quantum Rabi model (QRM) named as Rabi Rabi ; book describes the interaction of a two-level atom and a single-model optical field and can be viewed as the minimal version of the Dicke model Dicke . In the past years, there are efforts in finding the exactly analytically solutions of the QRM am_1 ; am_3 ; am_2 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 . Employing the Bargmann space representation bs , Braak obtained the transcendental function, namely the G-function, and found that the poles of the G-function were the exact solutions of QRM es1 . Quickly, Chen et al. obtained the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 . In the decade, the two methods have been used for calculating the exact solutions of the generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 . It was found that the symmetry breaking will lead to the topological transition in the QRM tp1 and the generalized QRM tp2 .
Accompanying by the search for the exact solutions, there are interests in studying the quantum criticality of the QRMs or the generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 . Accompanied by the quantum phase transition, i.e., the superradiance phase transition (SPT), there will occur macroscopic photon excitations R_d1 . In 2015, Hwang et al. studied the universal dynamics and the critical behaviors of the QRM and the exact phase boundary of SPT was obtained by an effective low-energy approximation HP . They found that the critical exponent of SPT $\nu$ is $\nu=\frac{3}{2}$ HP . By analyzing a general scaling function, Liu et al. found that the $\nu$ in the anisotropic QRM sill satisfies $\nu=\frac{3}{2}$ and the phase boundary of SPT is obtained by a semi-class method ps2 . Quickly, Wei et al. numerically verified the exact $\nu$ by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 . Subsequent researches show that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the crtical exponent is still equal to that of a single Rabi model, but for a three-level QRM, there occurs new critical exponent with $\nu=1$zhangyy_3 . We notice that the SPTs in the QRMs more or less depend on the atom-optical field coupling strength. It motivate us whether there exist an unconventional SPT which is independent of the atom-optical field coupling.
@@
Recently, the combination of the non-Hermiticity and the QRMs triggered intriguing physics. Cao et al. constructed a generalized non-Hermitian integrable QRM. They obtained the exact energy spectrum and eigenstates of this model by employing the Bethe ansatz es_NH_1 . The work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models. By introducing the parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, Lee et al. PT_Rabi_1 found that there was $\mathcal{PT}$ symmetry breaking PT ; PT1 accompanied by the real-complex transition in energy. With the exact solutions of the $PT$-symmetric Rabi models, one can locate the $PT$-symmetry breaking point, i.e., the exceptional point (EP) EP . Xie et al. investigated the Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model es_NH_2 . Under the multiple-photon resonance condition, they obtained the positions of the EPs from the exact Floquet spectrum. Employing the adiabatic approximation, Lu et al. obtained the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM, by which the EPs were located es_NH_3 . With the emergence of EP, observables present discontinuous behaviors. By the Bogoliubov transformation and G-function, Li et al. solved the exact energy spectra for a one-photon and two-photon $\mathcal{PT}$-symmetric QRM, and the zeros of G-function located the EPs es_NH_4 . These works provide an inspiration for us to study whether there will be a completely real energy spectrum in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In addition, we will study the quantum criticality of the model and discuss the similarity and difference with Hermitian cases.
@@
The paper is organized as follows. In Sec. II we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In Sec. III, we obtain the exact phase boundary of the model. In Sec. IV, we study the observables which signal the SPT. In Sec. V, we discuss the quantum criticality of the SPT. A experimental feasibility is discussed in Sec. VI, and a summary is presented in Sec. VII.
@@ II Two coupled non-Hermtian Rabi cavities
where $h_{j}$ describes non-Hermitian Rabi cavity $h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}$ with $j$ the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) the creation (annihilation) of photons. $\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field with $\lambda$ the weight of strength and $\varphi$ the coupling phase. The non-Hermiticity of the system requires $\varphi\in(-\pi,0)\cup(0,\pi)$.
@@
For a single non-Hermitian Rabi cavity, we argue that there is no parity-time ($\mathcal{PT}$) symmetry unless the coupling phase $\varphi=\pm\frac{\pi}{2}$. The time-reversal operator $\mathcal{T}$ dominates the complex conjugate operation, resulting in $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}$=$a^{{\dagger}}_{j}(a_{j})$. The parity operator $\mathcal{P}$ dominates the inversion operation, i.e., $\mathcal{P}\hat{x}_{j}\mathcal{P}$= $-\hat{x}_{j}$, and $\mathcal{P}\hat{p}_{j}\mathcal{P}$= $-\hat{p}_{j}$ PT . Thus, we have
@@
Only if $\varphi=\pm\frac{\pi}{2}$, the single non-Hermitian Rabi cavity is $\mathcal{PT}$-symmetric (see another prove in Ref. es_NH_4 ). For other values, it is not $\mathcal{PT}$-symmetric.
For the cascaded non-Hermitian QRM, we demonstrate that this system is completely not $\mathcal{PT}$-symmetric. $\mathcal{P}$ dominates a mirror operation between the two cavities, i.e., $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$, $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (here $j\neq j^{\prime}$), and $\mathcal{T}$ leads to $\mathcal{T}i\mathcal{T}=-i$. Performing $\mathcal{PT}$ on the cavity coupling term, we have
@@
Performing $\mathcal{PT}$ on $h_{j}$, we have
@@
$\neq h_{j^{\prime}}.$
Therefore, here the non-Hermitian cascaded QRM is not $\mathcal{PT}$-symmetric.
III Exact and unconventional superradiance phase transition
@@
As studied in Refs. ps1 ; ps2 , the SPT appears in the under the infinite $\eta$ limit. Under this condition, we can obtain the analytical phase boundary of SPT. When $\eta\rightarrow\infty$, $\mathcal{H}$ reduces to
$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x^{2}_{j}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
By diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$ as
$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}\pm\sqrt{1+2g^{2}x^{2}_{j}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
Performing the Taylor expansion on $E_{\pm}(x_{1},x_{2})$, and truncating them to the quadratic term $\mathcal{O}(x^{4}_{j})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (contains the ground-state information) becomes
@@
Evidently, the matrix $\Lambda$ consists of the real part ${\rm Re}(\Lambda)$ and the pure imaginary part ${\rm Im}(\Lambda)$ (which can be considered as the background loss), i.e., $\Lambda={\rm Re}(\Lambda)+{\rm Im}(\Lambda)$, with
@@
In fact, the phase boundary of SPT is nothing to do with background loss ${\rm Im}(\Lambda)$, but is determined by ${\rm Re}(\Lambda)$. After diagonalizing ${\rm Re}(\Lambda)$, the eigenvalues ϵ±subscriptitalic-$\epsilon_{\pm}$ of ${\rm Re}(\Lambda)$ is obtained as $\epsilon_{\pm}=1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)\pm 2\mathcal{J}$. When the low-energy branch in ϵ±subscriptitalic-$\epsilon_{\pm}$ equals to zero, i.e, $\epsilon_{-}=0$, we can extract the phase boundary of the superradiance phase transition, and it satisfies $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$. From the expression of the phase boundary, we can see that there are special phases, i.e., $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, at which the phase boundary is fixed at $\mathcal{J}=\frac{1}{2}$, and is independent of the atom-optical field coupling parameter $g$. Such type of SPT independent of atom-optical field coupling is counter-intuitive and has not been reported before, so we name this type of SPT as unconventional SPT. While $\gamma$ is taken at other values, the phase boundary depends on the $g$ and $\mathcal{J}$ together. Thus, we name this atom-optical field coupling dependent SPT as the conventional SPT. Without loss of generality, in the following, we will study the energy, photon population, qubit population, the fidelity and the fidelity susceptibility in the ground state at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the SPT of the non-Hermitian two-Rabi cavity system. The ground state here refers to the state with the smallest real part of the energy. Intuitively, the phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ are plotted in Fig. 1(a) and 1(b), respectively. The red regions denote the superradiance phases, separated from the normal phases (yellow regions) by the phase boundaries (blues curves).

### v2

# Phase driven unconventional superradiance phase transition in non-Hermitian cascaded quantum Rabi cavities

@@ Abstract
This study investigates phase-driven symmetry breaking leading to superradiance phase transitions in cascaded non-Hermitian quantum Rabi cavities. Non-Hermiticity is introduced via the phase coupling $\varphi$ between the atom and the optical field. In the thermodynamic limit of the quantum harmonic oscillator, we analytically derive the superradiance phase boundary, validated by observables. An unconventional quantum phase transition without a Hermitian analogue arises when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, where the phase boundary is uniquely determined by the cavity coupling, at $\mathcal{J}=\frac{1}{2}$, independent of the atom-photon coupling strength $g$. For other $\varphi$, the phase boundary relies on both $\mathcal{J}$ and $g$, similar to the scenario observed in Hermitian systems. Furthermore, we identify phase-driven first- and second-order superradiance phase transitions, focusing on the quantum criticality of the second-order transition by determining the critical exponents and the universality class. The feasibility of experimental realization is also discussed, aiming to inspire further studies on non-Hermitian superradiance quantum phase transitions.
@@ I Introduction
The cornerstone model for light-matter interaction is the quantum Rabi model (QRM), which involves a two-level atom (qubit) coupled through dipole interaction to a quantized single-mode cavity field and can be seen as the minimal version of the Dicke modelRabi ; book ; Dicke . In recent years, there have been efforts to find exact analytical solutions for the QRM am_1 ; am_2 ; am_3 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 . Using the Bargmann space representation bs , Braak obtained the transcendental function, known as the G-function, and discovered that the poles of the G-function were the exact solutions of the QRM es1 . Shortly after, Chen et al. achieved the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 . Over the past decade, these methods have been employed to calculate the exact solutions of generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 . It has been found that symmetry breaking can lead to topological transitions in both the QRM tp1 and generalized QRM tp2 .
Parallel to the search for exact solutions, significant interest has emerged in exploring the quantum criticality of QRMs and generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 . Quantum criticality denotes abrupt changes in a system’s properties at quantum phase transitions (QPT), where macroscopic observables display non-analytic behavior due to quantum fluctuations. In the context of light-matter interaction, this phenomenon is exemplified by quantum phase transitions such as the superradiance phase transition (SPT), characterized by macroscopic excitations in ground-state populationsreview .
Experimental observations in ion traps experiment have confirmed a second-order QPT in the QRM, where researchers noted a rapid increase in both phonon number and internal ion excitations at the critical point. In the deep strong coupling regime of light-matter interaction models, a notable feature is the zero-temperature photon vacuum instability driving systems to undergo a QPT at critical coupling strength $g_{c}$. Below $g_{c}$, the system resides in a normal symmetric phase with the field in vacuum and atoms in their lowest electronic state. Above $g_{c}$, both the field and atoms become excited, indicating a superradiant phase with broken symmetry. The thermodynamic limit is approached by increasing the number of particles to infinity in the Dicke model, resembling a classical limit for the atomic part. In contrast, phase transitions in single-atom systems occur at the classical oscillator limit obtained by increasing the qubit/oscillator frequency ratio.
Furthermore, the anisotropic quantum Rabi-Stark model (AQRSM), incorporating Stark coupling terms, demonstrates first-order and continuous QPTs xie2020first . Notably, excited-state quantum phase transitions (ESQPTs) have been extensively studied for both the standard quantum Rabi model (QRM) puebla2016excited and the AQRSM hu2023excited in the limit of infinite frequency ratio. These transitions involve non-analytic changes in the excited-state properties as a control parameter varies, occurring at finite energy densities and marked by singularities in the density of states. Studies have examined the universal dynamics and critical behaviors of the QRM, determining the exact phase boundary of the SPT using an effective low-energy approximation HP . The critical exponent of the SPT, $\nu$, was found to be $\nu=\frac{3}{2}$ HP . Analyzing a general scaling function has demonstrated that $\nu$ in the anisotropic QRM also satisfies $\nu=\frac{3}{2}$, and the phase boundary of the SPT was obtained via a semi-classical method ps2 . Subsequently, the exact $\nu$ was numerically confirmed by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 . Further research has revealed that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the critical exponent remains the same as that of a single Rabi model. However, for a three-level QRM, a new critical exponent with $\nu=1$ emerges zhangyy_3 . Notably, SPTs in QRMs depend on the atom-photon coupling strength, motivating investigation into the existence of unconventional SPTs independent of this coupling.
@@
Recently, the combination of non-Hermiticity and QRMs has sparked intriguing developments. A generalized non-Hermitian integrable QRM was constructed, and its exact energy spectrum and eigenstates were obtained using the Bethe ansatz es_NH_1 . This work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models. By introducing parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, it was discovered that $\mathcal{PT}$ symmetry breaking, accompanied by a real-complex transition in energy, occurs PT_Rabi_1 ; PT ; PT1 . With the exact solutions of the $\mathcal{PT}$-symmetric Rabi models, the $\mathcal{PT}$-symmetry breaking point, i.e., the exceptional point (EP), can be located EP . The Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model was investigated, and under multiple-photon resonance conditions, the positions of the EPs from the exact Floquet spectrum were determined es_NH_2 . Using the adiabatic approximation, the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM was obtained, locating the EPs es_NH_3 . With the emergence of EPs, observables present discontinuous behaviors. The exact energy spectra for one-photon and two-photon $\mathcal{PT}$-symmetric QRMs were solved using the Bogoliubov transformation and G-function, and the zeros of the G-function located the EPs es_NH_4 . Motivated by these studies, we will examine whether a fully real energy spectrum can emerge in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. Furthermore, this paper explores the quantum criticality of the model, revealing unique behaviors and transitions that distinguish it from conventional systems. By comparing and contrasting these findings with the established features of Hermitian systems, we aim to illuminate the intriguing parallels and divergences between the two realms.
@@
The paper is organized as follows. In Sec. II, we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In Sec. III, we obtain the exact phase boundary of the model. In Sec. IV, we study the observables which signal the SPT. In Sec. V, we discuss the quantum criticality of the SPT. Experimental feasibility is discussed in Sec. VI, and a summary is presented in Sec. VII.
@@ II Two coupled non-Hermtian Rabi cavities
where $h_{j}$ describes the non-Hermitian Rabi cavity
$h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x},$
with $j$ denoting the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) representing the creation (annihilation) operators of photons. Here, $\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field, where $\lambda$ represents the strength and $\varphi$ is the coupling phase. The non-Hermiticity of the system imposes that $\varphi\in(-\pi,0)\cup(0,\pi)$.
@@
In the context of a single non-Hermitian Rabi cavity, parity-time ($\mathcal{PT}$) symmetry is absent unless the coupling phase attains the specific values $\varphi=\pm\frac{\pi}{2}$. The time-reversal operator $\mathcal{T}$ enacts a complex conjugation operation, yielding $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}=a^{{\dagger}}_{j}(a_{j})$. Similarly, the parity operator $\mathcal{P}$ performs an inversion transformation defined by $\mathcal{P}\hat{x}_{j}\mathcal{P}=-\hat{x}_{j}$ and $\mathcal{P}\hat{p}_{j}\mathcal{P}=-\hat{p}_{j}$ PT . Consequently, we have
@@
Thus, the single non-Hermitian Rabi cavity manifests $\mathcal{PT}$-symmetry exclusively for $\varphi=\pm\frac{\pi}{2}$ (refer to an alternate proof in Ref. es_NH_4 ). For any other values of $\varphi$, $\mathcal{PT}$-symmetry is not exhibited.
For the cascaded non-Hermitian QRM, we verify that this system is entirely not $\mathcal{PT}$-symmetric. The parity operator $\mathcal{P}$ mirrors operations between the two cavities, specifically $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$ and $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (where $j\neq j^{\prime}$), while the time-reversal operator $\mathcal{T}$ satisfies $\mathcal{T}i\mathcal{T}=-i$. Performing $\mathcal{PT}$ on the cavity coupling term, we have
@@
Performing $\mathcal{PT}$ on $h_{j}$, we find that
@@
$\neq h_{j^{\prime}},$
where $h_{j}$ is the Hamiltonian of the $j$-th cavity. Therefore, the cascaded non-Hermitian QRM does not exhibit $\mathcal{PT}$-symmetry.
III Analytical Results
@@
As studied in Refs. ps1 ; ps2 , the superradiance phase transition (SPT) emerges under the infinite $\eta$ limit. Under this condition, we derive the analytical form of the Hamiltonian $\mathcal{H}$:
$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x_{j}^{2}+\sigma_{z}^{j}-\sqrt{2}g\sigma_{x}^{j}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
Diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$:
$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x_{j}^{2}\pm\sqrt{1+2g^{2}x_{j}^{2}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
Expanding $E_{\pm}(x_{1},x_{2})$ to quadratic terms $\mathcal{O}(x_{j}^{4})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (representing the ground-state information) simplifies to:
@@
The matrix $\Lambda$ separates into its real part ${\rm Re}(\Lambda)$ and pure imaginary part ${\rm Im}(\Lambda)$, representing the background loss:
@@
The phase boundary of the superradiance phase transition (SPT) is primarily determined by ${\rm Re}(\Lambda)$ rather than ${\rm Im}(\Lambda)$. Diagonalizing ${\rm Re}(\Lambda)$ yields the eigenvalues $\epsilon_{\pm}=1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)\pm 2\mathcal{J}$. The condition for the SPT phase boundary, where the low-energy branch $\epsilon_{-}=0$, results in $1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)=2\mathcal{J}$. This phase boundary reveals special phases at $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, where $\mathcal{J}=\frac{1}{2}$ independently of the atom-photon coupling parameter $g$. Such a phase transition, independent of the coupling $g$, is termed unconventional SPT as it has not been reported previously.
For other values of $\varphi$, the phase boundary depends jointly on $g$ and $\mathcal{J}$, categorizing it as conventional SPT. The subsequent analysis examines ground state energies, photon and qubit populations, fidelity, and fidelity susceptibility at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the non-Hermitian two-Rabi cavity system. Here, the ground state refers to the state with the lowest real part of the energy. Figures 1(a) and 1(b) illustrate phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, respectively, where red regions indicate superradiance phases separated by blue curves representing phase boundaries, and normal phases are depicted in yellow regions.